# SEA-SQL: Semantic-Enhanced Text-to-SQL with Adaptive Refinement

### v1

@@ 6 Conclusion
7 Limitations
This study presents several areas that necessitate further exploration and improvement. As shown in Appendix B, a significant portion of the errors generated by our methodology originated from the schema link. Consequently, future research may explore the effective schema linking techniques. Furthermore, while our approach imposes minimal overhead on the GPT-3.5 model, the prep-rocessing costs substantially increase when applied to the GPT-4 model. Therefore, the identification of strategies to maintain a reasonable overhead when employing the GPT-4 model comprises another pivotal area for future exploration.
Appendix A Results on Different Sample values
values num Simple Mod. Chall. All
zero 37.84 18.06 15.97 29.79
one 62.81 44.73 34.03 54.63
two 64.11 47.74 31.94 56.13
three 64.86 47.10 30.56 56.26
Table 6: The EX result of the number of different sample values on the BIRD development set.
We investigated the influence of different sample values on the results with the BIRD development set. Specifically, we tested with 0, 1, 2, and 3 samples as presented in Table 6. The results provided significant insights into the importance of sample values in the effectiveness of our methodology.
When the number of values is 0, essentially replacing our semantic-enhanced schema with the standard schema, a notable 56% decrease is observed. This reduction underscores the significance of our semantic-enhanced schema in achieving optimal results.
Comparatively, employing 2 or 3 sample values yields substantially superior results to using just 1 sample value. Although the performance with 3 sample values is marginally better than with 2, opting for 2 sample values notably enhances the model’s proficiency in generating SQL queries of moderate and challenging complexity levels. Considering the complexity of user queries, leveraging 2 sample values is more reasonable.
Appendix B Error Analysis
Figure 4: Error Distributions of SEA-SQL on the BIRD development set.
To better understand the constraints of our approach, we conducted an evaluation on the development set of the BIRD datasets. Following the methodology of DIN-SQL [Pourreza and Rafiei, 2023], we classified errors into five categories: schema linking errors, JOIN errors, GROUP BY errors, nesting errors, and other errors. Schema linking errors encompass the misuse of incorrect tables, columns, or values. JOIN errors are characterized by the use of incorrect columns or tables for JOIN operations. GROUP BY errors arise when incorrect columns are used or when GROUP BY is not utilized. Nesting errors involve the application of incorrect sub-queries or improper operator usage.

### v2

@@ 5.6 Robustness
5.7 LLM Backbones
To further verify the effectiveness of SEA-SQL, we conducted experiments across various LLM backbones. These experiments include tests on different generation models, such as GPT-4o-mini, Qwen-2.5-72b-it (Instruct) Yang et al. (2024), and Llama-3.1-70b-it Dubey et al. (2024), as well as various bias elimimnation models, like Gemma2-7B Rivière et al. (2024) and Llama2-7B Touvron et al. (2023). The results, presented in Table 7, clearly demonstrate that SEA-SQL consistently achieves excellent performance across all LLM backbones. Notably, SEA-SQL, when based on Llama-3.1-70b-it and Qwen-2.5-72b-it, even perform better than the ones with GPT-3.5 and GPT-4o-mini. However, to ensure a fair comparison, we chose GPT-3.5 as our primary experimental backbone.
5.8 Error Analysis
Figure 6: Error Distributions of SEA-SQL on the BIRD development set.
To better understand the constraints of our approach, we conducted an evaluation on the development set of the BIRD datasets. Following the methodology of DIN-SQL Pourreza and Rafiei (2023), we classified errors into five categories: schema linking errors, JOIN errors, GROUP BY errors, nesting errors, and other errors. Schema linking errors encompass the misuse of incorrect tables, columns, or values. JOIN errors are characterized by the use of incorrect columns or tables for JOIN operations. GROUP BY errors arise when incorrect columns are used or when GROUP BY is not utilized. Nesting errors involve the application of incorrect sub-queries or improper operator usage.
@@ 6 Conclusion
7 Acknowledgments
This work is supported by the National Natural Science Foundation of China (Nos. 62272054, 62192784,62172056), Beijing Nova Program (No. 20230484319), and Xiaomi Young Talents Program.
Appendix A Further Bias Elimination
After dynamic execution adjustment, new SQL quer-ies are generated by the LLM. To demonstrate that these queries do not introduce new biases, we conducted an experiment on the BIRD dataset by applying further bias elimination process after the dynamic execution adjustment. The results, shown in Table 8, indicate a decline in performance after this process. This decline is primarily because the dynamic execution adjustment maintains the original style of the SQL, ensuring it executable without introducing biases or syntax errors. Further bias elimination did not improve the results and might make the SQL non-executable.
Appendix B Few-shot Prompts
Previous work suggests that few-shot prompts may be more effective in leveraging LLMs for data-related tasks Zhang et al. (2024). Consequently, we explored whether using few-shot prompts would yield greater improvements, as shown in Table 9. The results indicate that using a 1-shot approach achieves better outcomes compared to zero-shot, and a 3-shot approach yields even higher results. However, the computational cost significantly increases with more examples. Given these costs, we continue to use zero-shot prompts for our main experiments. However, in scenarios where accuracy is a priority, using few-shot prompts tends to be more effective.